OpenStudy (anonymous):

Solve the equation 2x/x^2+2x-3+3/x+3=1 Please show steps this one got me the most baffled so far. Thanks in advance

hartnn (hartnn):

can you factorr x^2+2x-3 ??

hartnn (hartnn):

i assume its, \(\large \dfrac{2x}{x^2+2x-3}+\dfrac{3}{x+3}=1\)

OpenStudy (anonymous):

yes that is correct I was putting it in equation as we speak.

hartnn (hartnn):

so, factor, x^2+2x-3

OpenStudy (anonymous):

uhhh I need to check my notes again... I don't think I can

OpenStudy (anonymous):

oh wait just a sec

OpenStudy (anonymous):

x^2+x-3

hartnn (hartnn):

umm....what was that ? to factor x^2+2x-3, you need to find 2 numbers whose sum is +2 and product = -3 which are those numbers ?

OpenStudy (anonymous):

yup then I have no idea... would you be as kind as to show me how I would get those numbers? I have a bunch other I'll try tonight.

hartnn (hartnn):

" 2 numbers whose sum is +2 and product = -3" which numbers can have product as -3, very limited set....try to think, its easy

hartnn (hartnn):

like for example, if i need " 2 numbers whose sum is 5 and product = 6" it'd be 2 and 3

hartnn (hartnn):

it would be easier if you list out the factors of 3

hartnn (hartnn):

how about 3 and -1 ? required sum was 2

OpenStudy (anonymous):

thank you so much for guiding me through this btw...

OpenStudy (anonymous):

ok so now after I factor the numbers what next?

hartnn (hartnn):

x^2+2x-3 = (x+3)(x-1) right ?

OpenStudy (anonymous):

yes I believe so

hartnn (hartnn):

so we have \(\large \dfrac{2x}{(x+3)(x-1)}+\dfrac{3}{x+3}=1\)

hartnn (hartnn):

do you know what we can do, so as to make the denominator of both the terms same ?

OpenStudy (anonymous):

did we lost the original exponential x^2 under x2?

OpenStudy (anonymous):

oh nvm nvm nvm haha and yes we multiply (x-1) and get rid of it?

hartnn (hartnn):

or how about this ? \(\large \dfrac{2x}{(x+3)(x-1)}+\dfrac{3}{x+3}\times \dfrac{x-1}{x-1}=1\) multiplying and dividing by x-1 to make the denominators same ...

hartnn (hartnn):

did you notice what i did ?

OpenStudy (anonymous):

yes I think so in the first part you divided x-1 and on the second part multiply it, thus creating equal denominator

OpenStudy (anonymous):

2x/x-3+3/x-3-1 right?

OpenStudy (anonymous):

=1 I mean

hartnn (hartnn):

now since the denominator are same, we can combine the numerators, \(\large \dfrac{2x}{(x+3)(x-1)}+\dfrac{3}{x+3}\times \dfrac{x-1}{x-1}=1 \implies \large \dfrac{2x+3(x-1)}{(x+3)(x-1)}=1 \) ask if any doubts here.

hartnn (hartnn):

like \(\large \dfrac{☺}{♥}+\dfrac{☻}{♥}=\dfrac{☺+☻}{♥}\)

OpenStudy (anonymous):

yeah make it into simple fraction form

hartnn (hartnn):

now can you proceed further ? know the next step ?

OpenStudy (anonymous):

k let me try it

OpenStudy (anonymous):

I got 6x-3/x-3

OpenStudy (anonymous):

argh I feel I almost got it since x-3/x-3=1

hartnn (hartnn):

\( \large \dfrac{2x+3(x-1)}{(x+3)(x-1)}=1 \\ 2x+3(x-1)=(x+3)(x-1)\) just multiplied (x+3)(x-1) on both sides. now you'll get a quadratic in x

OpenStudy (anonymous):

give me a second I need to give my mother her meds be right back

hartnn (hartnn):

okay.

OpenStudy (anonymous):

thank you so much my friend I will have to recheck my notes regardless.

OpenStudy (anonymous):

as I am having trouble understanding it at all and I don't want a straight forward answer. All your help was much appreciated :)

hartnn (hartnn):

hmm.....do that and practice alot, you'll get it soon.